OpenStudy (anonymous):

18. Solve the inequality. Show your work. –6b > 42 or 4b > –4

OpenStudy (bibby):

for the first one, divide both sides by -6 since you're dividing by a negative, you'll have to flip the sign

OpenStudy (bibby):

for the second divide both sides by 4, no sign flipping needed

OpenStudy (anonymous):

Wait, do I divide 42 by -6?

OpenStudy (bibby):

you divide both sides, because you can't modify one side without doing the same thing to the other side for example if we had x+2>5 we'd subtract 2 from both sides and get \(x+2-2>5-2\\x\cancel{+2-2}>5-2\\x>3\)

OpenStudy (anonymous):

Oooh, so I'd divide -6 with 4b as well?

OpenStudy (anonymous):

Like this? \[42 \div -6 = -6 \div 4 ?\]

OpenStudy (bibby):

nah they're like 2 separate things, first we're solving \( –6b > 42\) for b then after we finish that, we'll solve \(4b > –4\) for b

OpenStudy (anonymous):

So b for the first one is -7 and b for the second one is -1?

OpenStudy (bibby):

these are inequalities, so b can be a lot of things. for example if we had x>2, x can be 3,55,1580157, or 2570123569723072893 so for the first one we have \(\dfrac{–6b}{-6} > \dfrac{42}{-6}\\b<-7\) and the 2nd one, we have \(4b > –4\\\dfrac{4b}{4} > \dfrac{-4}{4}\\b>-1\)

OpenStudy (anonymous):

Ok so, would is be: -6b > 42 or 4b > -4 -6 \[42 \div -6 = -7\] \[b = -7\] \[4 \div -4 = -1\] \[b = -1\]

OpenStudy (anonymous):

Thank you so much! You're a great help c:

OpenStudy (bibby):

you're very welcome c: please post more because I'm honestly not sure you fully understand the concept. you're converting inequalities to equalities

OpenStudy (anonymous):

I'm not supposed to do = I'm supposed to do < or >, correct?

OpenStudy (bibby):

yeah, and you keep it the same unless you're dividing by a negative, in which case you flip it you'll probably find better help on youtube as far as learning what to do, when to do, and why to do it

OpenStudy (anonymous):

Ok, thanks so much again! c:

OpenStudy (bibby):

you can tag me in the future using " @bibby " no quotes

OpenStudy (anonymous):

Okay (: